# HeMiTo-dynamics: a characterisation of mammalian prion toxicity using non-dimensionalisation, linear stability and perturbation analyses

### v1

@@
Prion-like proteins play crucial parts in biological processes in organisms ranging from yeast to humans. For instance, many neurodegenerative diseases are believed to be caused by the production of prion-like proteins in neural tissue. As such, understanding the dynamics of prion-like protein production is a vital step toward treating neurodegenerative disease. Mathematical models of prion-like protein dynamics show great promise as a tool for predicting disease trajectories and devising better treatment strategies for prion-related diseases. Herein, we investigate a generic model for prion-like dynamics consisting of a class of non-linear ordinary differential equations (ODEs), establishing constraints through a linear stability analysis that enforce the expected properties of mammalian prion-like toxicity. Furthermore, we identify that prion toxicity evolves through three distinct phases for which we provide analytical descriptions using perturbation analyses. Specifically, prion-toxicity is initially characterised by the healthy phase, where the dynamics are dominated by the healthy form of prions, thereafter the system enters the mixed phase, where both healthy and toxic prions interact, and lastly, the system enters the toxic phase, where toxic prions dominate, and we refer to these phases as HeMiTo-dynamics. These findings hold the potential to aid researchers in developing precise mathematical models for prion-like dynamics, enabling them to better understand underlying mechanisms and devise effective treatments for prion-related diseases.
Keywords:
Prions, non-linear ordinary differential equations (ODEs), non-dimensionalisation, linear stability analysis, perturbation analysis.
@@ 1 Introduction
Prions are a class of proteins that are responsible for diseases such as Creutzfeldt–Jakob disease and transmissible spongiform encephalitis [4, 12]. Prions typically consist of several structural variants, some of which are healthy and some of which are harmful. The harmful, also called misfolded, prion variants can bind to healthy variants, misfolding them, and effectively converting them into toxic, misfolded protein. Misfolded prion proteins also form aggregates which eventually induce cell death. Moreover, many other proteins have been shown to have prion-like features, such as having a toxic variant of itself capable of converting healthy protein into toxic protein. These prion-like proteins are believed to underlie neurodegenerative diseases in mammals [3, 1]. A notable example is that of amyloid beta playing a critical role in Alzheimer’s disease [2, 3]. Longitudinal studies tracking biomarkers of Alzheimer’s disease over time indicate that the toxic form increases like a sigmoid curve over time [7]. However, due to the long time scales of human neurodegeneration, it is difficult to investigate the biological mechanisms underlying prion dynamics in human brains. To this end, prions are also studied in fungi in general and baker’s yeast in particular [14] where prion toxicity occurs on much shorter time scales. A powerful tool for inferring biological mechanisms underlying prion dynamics is mechanistic modelling, and hitherto, numerous mathematical models based on the prion hypothesis have been constructed and analysed [18, 13, 16, 15, 10, 6] which, for instance, capture the sigmoid-like accumulation of toxic prions over time.
@@
There are different levels of detail that can be used to describe the process by which healthy proteins are transformed into toxic proteins. When proteins become toxic, they form aggregates of various sizes, which can be modelled explicitly [18, 6], describing the concentration evolution of each aggregate size. However, these models of aggregation are infinite-dimensional and depend on a large number of parameters, which makes numerical simulations and data fitting quite challenging. The heterodimer model describes a simplification of prion-like replication, where aggregates are ignored. Instead, one only considers a concentration of healthy proteins and a concentration of toxic proteins, for which a reaction is transforming healthy into toxic protein [14, 18]. In the context of prion dynamics in the baker’s yeast Saccharomyces cerevisiae, a class of heterodimer models encapsulating other previously analysed models of mammalian prion toxicity have been recently proposed by Lemarre et al. [9]. Denoting the concentration of the healthy form by $H(t)$ and the concentration of the toxic form by $T(t)$ where the independent variable $t$ corresponds to time, we consider a heterodimer model with an arbitrary conversion function $f$ given by
@@
where time derivatives are denoted by dots, e.g. $\dot{H}=\mathrm{d}H/\mathrm{d}t$. Here, we assume that the concentrations of the two species $H$ and $T$ are measured in [nM], that time $t$ is measured in [years], and that the initial conditions in Eq. (3) are defined by positive constants $H_{0},T_{0}>0$. Then $k_{1}\,[\mathrm{nM}/\mathrm{year}]$ is the constant formation rate of the healthy species $H$, $k_{2}$ and $k_{4}$ are first order degradation rates of $H$ and $T$ measured in $[\mathrm{year}^{-1}]$, $k_{3}\,[\mathrm{nM}^{-1}\mathrm{year}^{-1}]$ is the second order conversion rate from the healthy to the toxic form and $f(T)$ is the conversion function which is a unitless bounded analytical function affecting the conversion rate. Mathematically speaking, both in the context of fungi and mammals, the choice of the conversion function $f$ must support the existence of two steady states, namely a so-called healthy steady state (HSS) which is free of toxic prions and a so called toxic steady state (TSS) characterised by toxic prions. Given the existence of two such steady states, the desired dynamical properties of prion models differ slightly for mammals and fungi.
@@
When it comes to prion dynamics in Saccharomyces cerevisiae, a desired mathematical property is that of bistability [9]. Here, toxic prions are inherited by daughter cells from mother cells after cell division, and it is the initial concentration of toxic prions that determines whether the daughter cell will be functional or dysfunctional. Mathematically, this feature is captured by bistability implying that both the HSS and the TSS are stable simultaneously. In this case, it is the amount of toxic prions that the daughter inherits after cell division corresponding to $T_{0}$ in Eq. (3) that determines whether the system evolves to the HSS or the TSS. A concrete example of a bistable model investigated by Lemarre et al. [9] is characterised by choosing the conversion function to a Hill function of the type $f(T)=T^{n-1}/(K+T^{n})$ for some positive integer $n$ and some positive constant $K$.
@@
For prion models of human and mammal neurodegeneration, the concentration of the toxic form of prions is initially low and thereafter it increases over time like a sigmoid curve [7]. Importantly, since the initial concentration of the toxic form $T_{0}$ in Eq. (3) is low, conditions on the parameters and the conversion function $f$ are typically enforced in order to capture the sigmoidal accumulation of the toxic form. Technically, these conditions result in stability critiera such that the HSS is a saddle point and the TSS is a stable node. A concrete example of such a prion model describing the dynamics of the biomarkers amyloid beta and tau proteins in the context of Alzheimer’s disease was proposed by Thompson et al. [16] and it corresponds to the conversion function $f(T)=1$.
@@
Nevertheless, given this latter prion model of human neurodegeneration, there are two fundamental unanswered questions. What choices of conversion function characterise models of mammalian prion toxicity where the HSS is a saddle point and the TSS is a stable node? Also, these stability criteria are ensured by means of linear stability analysis yielding knowledge about the long term dynamics of the models at hand. However, another unknown question is what features characterise the early dynamics of this type of models starting from low initial concentrations $H_{0}$ and $T_{0}$ in Eq. (3)? Critically, a linear stability analysis cannot provide such detailed descriptions of the early dynamics, and typically detailed knowledge about dynamical properties of ODEs at different time scales can be obtained by means of a perturbation analysis [5]. Given an appropriate non-dimensionalisation of the model of interest resulting in a small perturbation parameter $\varepsilon\ll{1}$, solutions are expressed as series expansions in this perturbation parameter where the different orders in the series expansions approximate the dynamics of the system of interest at certain time scales. In particular, the initial dynamics is captured by so called outer solutions corresponding to the $\mathcal{O}(1)$ terms in the perturbation series, and typically the original system of ODEs can be expressed as simpler ODEs that are valid on specific time scales encoded by orders of the perturbation parameter $\varepsilon$.
@@
In this work, we derive a class of prion models of mammalian prion toxicity. Using a non-dimensionalisation resulting in a class of models defined by a perturbation parameter corresponding to a dimensionless conversion rate, we define conditions on the conversion function $f$ such that the HSS is a saddle point and the TSS is a stable node. Using numerical simulations starting from low initial concentrations of prions, we show that the dynamics of prion models in this class are divided into three phases referred to as HeMiTo-dynamics. First, the system goes through the healthy phase called the He-phase where the concentration of the healthy form increases and while that of the toxic form is constant. Thereafter, the system goes through the mixed phase called the Mi-phase where the healthy form reaches a maximum value and the toxic form increases. Lastly, the system goes through the toxic phase called the To-phase where the toxic form dominates the dynamics and specifically the system evolves towards the TSS. Using perturbation analyses, we derive approximations of analytical solutions in the initial He- and Mi-phases capturing the exact time dependence of solutions of our class of prion models and we validate these approximations by fitting them to numerical solutions. Lastly, we show that the choice of conversion function $f$ results in two main types of dynamics during the Mi-phase where the concentration profile for the healthy form $u(\tau)$ is either a concave function which reaches a clear maximum value similar to the epidemiological SIR model [8] or it behaves more like a logistic growth function evolving towards a carrying capacity.
@@
In order for a TSS to exist, we require that both $H_{2}^{\star}$ and $T_{2}^{\star}$ are positive. Accordingly, the conversion function must satisfy $f(T_{2}^{\star})>0$ so that $H_{2}^{\star}>0$ and it must also be chosen so that $\varepsilon$ in Eq. (7) lies in the interval $\varepsilon\in(0,1)$ so that $T_{2}^{\star}>0$. Consequently, $\varepsilon$ is our perturbation parameter which we subsequently use as the basis for our non-dimensionalisation and perturbation analysis. In light of this perturbation parameter, we introduce the following dimensionless time variable
@@ 3 Defining a class of models of mammalian prion toxicity by means of linear stability analysis
In terms of the function $g$ in Eq. (17), the inequality in Eq. (18) determining when the HSS is a saddle point is given by $f(0)>g(0)$. This fact together with Remark 17 implies that the function $f$ is bounded from below by the function $g$ which is expressed as follows
@@ 4 HeMiTo-dynamics: prion toxicity is characterised by three phases
The dynamics characterising mammalian prion toxicity rely on two biologically imposed conditions. First, the initial concentrations of both forms of prions are low and second the concentration of the toxic form increases and accumulates over time. Mathematically, these assumptions correspond to low initial concentrations $u_{0}=\mathcal{O}(1)<\mathcal{O}(1/\varepsilon)$ and $v_{0}=\mathcal{O}(1)<\mathcal{O}(1/\varepsilon)$ in Eq. (13) and that parameters and conversion functions $f$ are chosen such that the HSS is a saddle point and the TSS is a stable node in accordance with Theorem 2. Additionally, we assume that the conversion function $f(v)$ is of the same order as $v$, and therefore we have $f(v_{0})=\mathcal{O}(1)$ initially. Given these conditions, the dynamics characterising mammalian prion toxicity are divided into three phases captured by the acronym HeMiTo (FIG. 2). First, the system goes through the healthy phase referred to as the He-phase where the healthy form dominates the dynamics meaning that its concentration increases while the concentration of the toxic form is constant. Given that the HSS is a saddle point and that the initial conditions are chosen relatively close to the HSS in the $(u,v)$ phase plane, solution trajectories in the He-phase move along the stable eigenvector of the linearised system around the HSS. Second, the system goes through the mixed phase referred to as the Mi-phase where both species interact. Specifically, the trajectories of the healthy form attain maxima while the trajectories of the toxic form increase over time during the Mi-phase. Third, the system goes through the toxic phase referred to as the To-phase where the toxic form dominates the dynamics. Mathematically, the dynamics of the system in the To-phase are approximately captured by the linearised system around the TSS which was previously considered in the linear stability analysis. Subsequently, we present similar approximations describing the dynamics during the He- and Mi-phases.
@@
Figure 2: HeMiTo-dynamics. The dynamics of the system of ODEs in Eqs. (11) and (12) characterising mammalian prion toxicity are divided into three phases remembered by the acronym HeMiTo. First, the system goes through the healthy phase referred to as the He-phase where the healthy form $u(\tau)$ increases and the toxic form $v(\tau)$ is constant. Second, the system goes through the mixed phase referred to as the Mi-phase where the healthy and toxic forms interact. Specifically, the healthy form reaches a maximum value while the toxic form increases throughout the Mi-phase. Third, the system goes through the toxic phase referred to as the To-phase where the toxic form dominates the dynamics over the healthy form. Here, solutions of the class of prion models evolve towards the stable TSS. This type of dynamics is visualised in two cases defined by distinct conversion functions: (A) $f(v)=1.00$, and (B) $f(v)=0.57+(v^{4}/(1+v^{5}))$. The parameters defining the illustrated curves are $c_{1}=1.75$, $c_{2}=0.70$ and $\varepsilon=0.40$, and the initial conditions in both cases are given by $(u_{0},v_{0})=(0.20,0.05)$.
@@ 5 Finding approximate analytical solutions in the He-phase by means of perturbation ansätze
Initially, the dynamics are dominated by the formation of intact prions while the conversion from healthy to toxic prions is comparatively small. Provided low initial prion concentrations in Eq. (13) implying that $u_{0}=\mathcal{O}(1)$ and $v_{0}=\mathcal{O}(1)$, this means that the formation rate “$c_{1}$” and the degradation rate “$c_{2}u$” in the ODE for $u$ in Eq. (11) are of the same order while the conversion rate “$\varepsilon{u}vf(v)$” is small, and in particular we have $u(\tau)=\mathcal{O}(1)$ for early times $\tau$ close to 00. Accordingly, consider the following regular perturbation ansätze for the healthy and toxic form, respectively:
@@
During the Mi-phase, the conversion rate from healthy to toxic prions dominates the dynamics over the formation rate of the healthy form. In particular, since the He-phase results in a build-up of the healthy form, we have that $u(\tau)=\mathcal{O}(1/\varepsilon)$ which implies that the constant formation rate $c_{1}=\mathcal{O}(1)<\mathcal{O}(1/\varepsilon)$ is comparatively small during the Mi-phase. To capture the dynamics in this phase, we re-scale the states as follows
@@
where $v_{0}=\mathcal{O}(1)$ is the initial concentration of the toxic form in Eq. (13) and where $f(v_{0})=\mathcal{O}(1)$. This approximation can be justified by the arguments presented by Gerlee [5] which state that the approximation is, in fact, exact when the conversion function $f$ is a polynomial. Also, the approximation is accurate for non-polynomial conversion functions $f$ as they, in turn, can be approximated by polynomials. In light of this approximation for the conversion function $f$ in Eq. (27), by multiplying the original ODEs for $u$ and $v$ in Eqs. (11) and (12) by $\varepsilon$, we obtain an approximate system of ODEs describing the dynamics during the Mi-phase in terms of the new states $U$ and $V$ in Eq. (26) which is given by
@@
Given this system, we find asymptotic approximations of the solutions $u_{\mathrm{Mi}}(\tau)$ and $v_{\mathrm{Mi}}(\tau)$ that explicitly describe the time dependence. Importantly, the structure of this model of prion dynamics during the Mi-phase is similar to the well-known epidemiological SIR model originally formulated by McKendrick and Kermack [8]. Recently asymptotic approximations of the solutions of the SIR model was found by integrating the ODEs of interest with respect to time and thereafter approximating the resulting unknown integrals [11]. Using this technique, the ODEs for $u_{\mathrm{Mi}}$ and $v_{\mathrm{Mi}}$ in Eqs. (32) and (33), respectively, can be formulated as equivalent integral equations
$u_{\mathrm{Mi}}(\tau)$ $=\tilde{C}_{1}\exp\left(-c_{2}\tau\right)\exp\left(-f(v_{0})\int_{0}^{\tau}v(s)\mathrm{d}s\right)\,,$ (34)

### v2

@@ 1 Introduction
Prions are a class of proteins that are responsible for diseases such as Creutzfeldt-Jakob disease, kuru, and bovine spongiform encephalopathy [5, 13]. Prions typically consist of several structural variants, some of which are healthy and some of which are harmful. The harmful, also called misfolded, prion variants can bind to healthy variants, misfolding them, and effectively converting them into toxic, misfolded protein. Misfolded prion proteins also form aggregates which eventually induce cell death. Moreover, many other proteins have been shown to have prion-like features such as having a toxic variant of itself capable of converting healthy protein into toxic protein. These prion-like proteins are believed to underlie neurodegenerative diseases in mammals [4, 1]. A notable example is that of amyloid beta playing a critical role in Alzheimer’s disease [3, 4]. Longitudinal studies tracking biomarkers of Alzheimer’s disease over time indicate that the toxic form increases like a sigmoid curve over time [8]. However, due to the long time scales of human neurodegeneration, it is difficult to investigate the biological mechanisms underlying prion dynamics in human brains. Prion-like proteins are also studied in fungi in general and baker’s yeast in particular [15], where such proteins often participate in cellular processes beneficial to the organism [2]. Prion-like protein dynamics in yeast operate on much shorter time scales and may thus be studied in greater detail. A powerful tool for inferring biological mechanisms underlying prion dynamics is mechanistic modelling, and hitherto, numerous mathematical models based on the prion hypothesis have been constructed and analysed [20, 14, 18, 17, 11, 7] which, for instance, capture the sigmoid-like accumulation of toxic prions over time.
@@
Prion-like dynamics can be described in different levels of detail. Note that we will refer to the non-aggregating protein variant as healthy and the aggregating variant as toxic. However, keep in mind that the aggregating variant could potentially be beneficial as opposed to pathological, in particular when discussing fungal prions. When healthy prion-like proteins are transformed into their toxic counterpart, they form aggregates of various sizes, which can be modelled explicitly [20, 7], describing the concentration evolution of each aggregate size. However, these models of aggregation are infinite-dimensional and depend on a large number of parameters, which makes numerical simulations and data fitting quite challenging. The heterodimer model describes a simplification of prion-like replication, where aggregates are ignored. Instead, one only considers a concentration of healthy proteins and a concentration of toxic proteins, for which a reaction is transforming healthy into toxic protein [15, 20]. In the context of prion dynamics in the baker’s yeast Saccharomyces cerevisiae, a class of heterodimer models encapsulating other previously analysed models of mammalian prion toxicity have been recently proposed by Lemarre et al. [10]. Denoting the concentration of the healthy form by $H(t)$ and the concentration of the toxic form by $T(t)$ where the independent variable $t$ corresponds to time, we consider a heterodimer model with an arbitrary conversion function $f$ given by
@@
where time derivatives are denoted by dots, e.g. $\dot{H}=\mathrm{d}H/\mathrm{d}t$. Here, we assume that the concentrations of the two species $H$ and $T$ are measured in [nM], that time $t$ is measured in [years], and that the initial conditions in Eq. (3) are defined by positive constants $H_{0},T_{0}>0$. Then $k_{1}\,[\mathrm{nM}/\mathrm{year}]$ is the constant formation rate of the healthy species $H$, $k_{2}$ and $k_{4}$ are first order degradation rates of $H$ and $T$ measured in $[\mathrm{year}^{-1}]$, $k_{3}\,[\mathrm{nM}^{-1}\mathrm{year}^{-1}]$ is the second order conversion rate from the healthy to the toxic form and $f(T)$ is the conversion function which is a unitless bounded analytical differentiable function affecting the conversion rate. Mathematically speaking, both in the context of fungi and mammals, the choice of the conversion function $f$ must support the existence of two steady states, namely a so-called healthy steady state (HSS) which is free of toxic prions and a so called toxic steady state (TSS) characterised by toxic prions. Given the existence of two such steady states, the desired dynamical properties of prion models differ slightly for mammals and fungi.
@@
When it comes to prion dynamics in Saccharomyces cerevisiae, a desired mathematical property is that of bistability [10]. Here, prions are inherited by daughter cells from mother cells after cell division, and it is the initial concentration of prion aggregates that determines whether the daughter cell will be functional or dysfunctional. Mathematically, this feature is captured by bistability implying that both the HSS and the TSS are stable simultaneously. In this case, it is the amount of prion aggregates that the daughter inherits after cell division corresponding to $T_{0}$ in Eq. (3) that determines whether the system evolves to the HSS or the TSS. A concrete example of a bistable model investigated by Lemarre et al. [10] is characterised by choosing the conversion function to a Hill function of the type $f(T)=K_{1}T^{n-1}/(K_{2}+T^{n})$ for some positive integer $n$ and some positive constants $K_{1},K_{2}$.
@@
For prion models of human and mammal neurodegeneration, the concentration of the toxic form of prions is initially low and thereafter it increases over time like a sigmoid curve [8]. Importantly, since the initial concentration of the toxic form $T_{0}$ in Eq. (3) is low, conditions on the parameters and the conversion function $f$ are typically enforced in order to capture the sigmoidal accumulation of the toxic form. Technically, these conditions result in stability critiera such that the HSS is a saddle point and the TSS is a stable node. A concrete example of such a prion model describing the dynamics of the biomarkers amyloid beta and tau proteins in the context of Alzheimer’s disease was proposed by Thompson et al. [18] and it corresponds to the conversion function $f(T)=1$. Note that the initial concentrations of toxic prions in mammals are not a result of skewed protein distributions during cell replication, but rather variations in the propensity of brain regions to form prion-like pathology.
@@
Nevertheless, given this latter prion model of human neurodegeneration, there are two fundamental unanswered questions. What choices of conversion function characterise models of mammalian prion toxicity where the HSS is a saddle point and the TSS is a stable node? Also, these stability criteria are ensured by means of linear stability analysis yielding knowledge about the long term dynamics of the models at hand. However, another unknown question is what features characterise the early dynamics of this type of model starting from low initial concentrations $H_{0}$ and $T_{0}$ in Eq. (3)? Critically, a linear stability analysis cannot provide such detailed descriptions of the early dynamics, and typically detailed knowledge about dynamical properties of ODEs at different time scales can be obtained by means of a perturbation analysis [6]. Given an appropriate non-dimensionalisation of the model of interest resulting in a small perturbation parameter $\varepsilon\ll{1}$, solutions are expressed as series expansions in this perturbation parameter where the different orders in the series expansions approximate the dynamics of the system of interest at certain time scales. In particular, the initial dynamics are captured by so called outer solutions corresponding to the $\mathcal{O}(1)$ terms in the perturbation series, and typically the original system of ODEs can be expressed as simpler ODEs that are valid on specific time scales encoded by orders of the perturbation parameter $\varepsilon$.
@@
In this work, we derive a class of prion models of mammalian prion toxicity. Using a non-dimensionalisation resulting in a class of models defined by a perturbation parameter corresponding to a dimensionless conversion rate, we define conditions on the conversion function $f$ such that the HSS is a saddle point and the TSS is a stable node. Using numerical simulations starting from low initial concentrations of prions, we show that the dynamics of prion models in this class are divided into three phases referred to as HeMiTo-dynamics. First, the system goes through the healthy phase called the He-phase, where the concentration of the healthy form normalises while that of the toxic form does not change. Thereafter, the system goes through the mixed phase called the Mi-phase where the healthy form reaches a maximum value and the toxic form increases. Lastly, the system goes through the toxic phase called the To-phase where the toxic form dominates the dynamics and specifically the system evolves towards the TSS. In many scenarios, it would be reasonable to assume that the healthy protein form has already normalised to regular physiological levels. In this case, the He-phase has already passed, and the introduction of a seed or alteration in system parameters introduces the Me-phase. Moreover, using perturbation analyses, we derive approximations of analytical solutions in the initial He- and Mi-phases capturing the exact time dependence of solutions of our class of prion models and we validate these approximations by fitting them to numerical solutions. Lastly, we show that the choice of conversion function $f$ results in two main types of dynamics during the Mi-phase where the concentration profile for the healthy form $u(\tau)$ is either a concave function which reaches a clear maximum value similar to the epidemiological SIR model [9] or it behaves more like a logistic growth function evolving towards a carrying capacity.
@@
In order for a biologically-reasonable TSS to exist, we require that both $H_{2}^{\star}$ and $T_{2}^{\star}$ are positive. Accordingly, the conversion function must satisfy $f(T_{2}^{\star})>0$ so that $H_{2}^{\star}>0$, and $f(T_{2}^{\star})$ must be chosen so that $\varepsilon$ in Eq. (7) lies in the interval $\varepsilon\in(0,1)$ which ensures that $T_{2}^{\star}>0$ according to Eq. 6. Consequently, $\varepsilon$ is our perturbation parameter which we subsequently use as the basis for our non-dimensionalisation and perturbation analysis. In light of this perturbation parameter, we introduce the following dimensionless time variable
@@ 3 Defining a class of models of mammalian prion toxicity by means of linear stability analysis
In terms of the function $g$ in Eq. (17), the inequality in Eq. (18) determining when the HSS is a saddle point is given by $f(0)>g(0)$. This fact together with Remark 17 implies that the function $f$ is bounded from below by the function $g$ until their first intersection, which is expressed as follows
@@ 4 HeMiTo-dynamics: prion toxicity is characterised by three phases
The dynamics of mammalian prion toxicity characterised herein rely on two imposed conditions. First, the initial concentrations of healthy and toxic proteins are low, and second, the concentration of the toxic form increases and accumulates over time. Mathematically, these assumptions correspond to low initial concentrations $u_{0}=\mathcal{O}(1)<\mathcal{O}(1/\varepsilon)$ and $v_{0}=\mathcal{O}(1)<\mathcal{O}(1/\varepsilon)$ in Eq. (13) and that parameters and conversion functions $f$ are chosen such that the HSS is a saddle point and the TSS is a stable node in accordance with Theorem 2. Additionally, we assume that the conversion function $f(v)$ and toxic concentration $v$ is of the same order $f(v)\sim\mathcal{O}(v)$, and therefore we have $f(v_{0})=\mathcal{O}(1)$ initially. Given these conditions, the dynamics characterising mammalian prion toxicity are divided into three phases captured by the acronym HeMiTo (FIG. 2). First, the system goes through the healthy phase referred to as the He-phase where the healthy form dominates the dynamics. In this phase, the healthy concentration approaches the HSS, and the concentration of the toxic form is constant (and low). Given that the HSS is a saddle point and that the initial conditions are chosen relatively close to the HSS in the $(u,v)$ phase plane, solution trajectories in the He-phase move along the stable eigenvector of the linearised system around the HSS. Second, the system goes through the mixed phase referred to as the Mi-phase where both species interact. Specifically, the trajectories of the healthy form attain maxima while the trajectories of the toxic form increase over time during the Mi-phase. Third, the system goes through the toxic phase referred to as the To-phase where the toxic form dominates the dynamics. Mathematically, the dynamics of the system in the To-phase are approximately captured by the linearised system around the TSS which was previously considered in the linear stability analysis. Subsequently, we present similar approximations describing the dynamics during the He- and Mi-phases. It is, however, important to note that in biological settings, the initial concentration of healthy protein would not be small, but rather, close to the HSS. In this case, the healthy phase is trivial and characterised by constant concentrations of the healthy form and low concentrations of the toxic form. This point is addressed further in the Discussion.
@@
Figure 2: HeMiTo-dynamics. The dynamics of the system of ODEs in Eqs. (11) and (12) characterising mammalian prion toxicity are divided into three phases remembered by the acronym HeMiTo. First, the system goes through the healthy phase referred to as the He-phase where the healthy form $u(\tau)$ approaches the healthy steady-state, and the toxic form $v(\tau)$ is constant. Second, the system goes through the mixed phase referred to as the Mi-phase where the healthy and toxic forms interact. Specifically, the healthy form begins to decrease while the toxic form increases throughout the Mi-phase. Third, the system goes through the toxic phase referred to as the To-phase where the toxic form dominates the dynamics over the healthy form. Here, solutions of the class of prion models evolve towards the stable TSS. This type of dynamics is visualised in two cases defined by distinct conversion functions: (A) $f(v)=1.00$, and (B) $f(v)=0.57+(v^{4}/(1+v^{5}))$. The parameters defining the illustrated curves are $c_{1}=1.75$, $c_{2}=0.70$ and $\varepsilon=0.40$, and the initial conditions in both cases are given by $(u_{0},v_{0})=(0.20,0.05)$.
@@ 5 Finding approximate analytical solutions in the He-phase by means of perturbation ansätze
Initially, the dynamics are dominated by the formation of healthy prions while the conversion from healthy to toxic prions is comparatively small. Provided low initial prion concentrations in Eq. (13) implying that $u_{0}=\mathcal{O}(1)$ and $v_{0}=\mathcal{O}(1)$, this means that the formation rate “$c_{1}$” and the degradation rate “$c_{2}u$” in the ODE for $u$ in Eq. (11) are of the same order while the conversion rate “$\varepsilon{u}vf(v)$” is small, and in particular we have $u(\tau)=\mathcal{O}(1)$ for early times $\tau$ close to 00. Accordingly, consider the following regular perturbation ansätze for the healthy and toxic form, respectively:
@@
During the Mi-phase, the conversion rate from healthy to toxic prions dominates the dynamics over the formation rate of the healthy form. In particular, since the healthy form approaches the HSS in the He-phase, we have that $u(\tau)=\mathcal{O}(1/\varepsilon)$ which implies that the constant formation rate $c_{1}=\mathcal{O}(1)<\mathcal{O}(1/\varepsilon)$ is comparatively small during the Mi-phase.
To capture the dynamics in this phase, we re-scale the states as follows
@@
where $v_{0}=\mathcal{O}(1)$ is the initial concentration of the toxic form in Eq. (13) and where $f(v_{0})=\mathcal{O}(1)$. This approximation can be justified by the arguments presented by Gerlee [6] which state that the approximation is, in fact, exact when the conversion function $f$ is a polynomial. Also, the approximation is accurate for non-polynomial conversion functions $f$ granted continuity over a closed interval, as they, in turn, can be approximated by polynomials as guaranteed by the Stone-Weierstrass theorem [16]. In light of this approximation for the conversion function $f$ in Eq. (27), by multiplying the original ODEs for $u$ and $v$ in Eqs. (11) and (12) by $\varepsilon$, we obtain an approximate system of ODEs describing the dynamics during the Mi-phase in terms of the new states $U$ and $V$ in Eq. (26) which is given by
@@
Given this system, we find asymptotic approximations of the solutions $u_{\mathrm{Mi}}(\tau)$ and $v_{\mathrm{Mi}}(\tau)$ that explicitly describe the time dependence. Importantly, the structure of this model of prion dynamics during the Mi-phase is similar to the well-known epidemiological SIR model originally formulated by McKendrick and Kermack [9]. Recently, asymptotic approximations of the solutions of the SIR model were found by integrating the ODEs of interest with respect to time and thereafter approximating the resulting unknown integrals [12]. Using this technique, the ODEs for $u_{\mathrm{Mi}}$ and $v_{\mathrm{Mi}}$ in Eqs. (32) and (33), respectively, can be formulated as equivalent integral equations
$u_{\mathrm{Mi}}(\tau)$ $=\tilde{C}_{1}\exp\left(-c_{2}\tau\right)\exp\left(-f(v_{0})\int_{0}^{\tau}v_{\mathrm{Mi}}(s)\mathrm{d}s\right)\,,$ (34)